OpenStudy (anonymous):

(6c) ^9 (6c)^ 5 simplify

OpenStudy (callisto):

\[(6c) ^9 (6c)^ 5\]\[= (6c)^{9+5} = (6c)^{14} = 6^{14}c^{14}\] Can you do the last step? simplify 6^ (14)

OpenStudy (anonymous):

not sure how

OpenStudy (callisto):

Hmm.. Do you understand the previous steps?

OpenStudy (anonymous):

ya

OpenStudy (callisto):

6^14 = 6 x 6 x 6x 6x 6x 6x 6x 6x 6x 6x 6x 6x 6x 6 =?